OpenStudy (anonymous):

Show that area of Pythagorean triangle x,y,z cannot be square (U can use that a^4-b^4 = c^2 has no positive solution)

OpenStudy (anonymous):

the base and height of pythagorean triangle are 3d and 4d where d is a constant Area = 1/2 x 3d x 4d = 3d x 2d = 6d^2 this is never a square

OpenStudy (anonymous):

Its not generalised though!!

OpenStudy (anonymous):

It's a start... Let's say we have x^2 + y^2 = z^2 and xy = 2n^2 (triangle area)

OpenStudy (anonymous):

Okay so we take it as hypothesis and prove it a contradiction..I'll try

OpenStudy (experimentx):